how many moles of HNO3 will be produced from the reaction of 63.5g of NO2 with excess water in the following chemical reaction 3NO2 + H2O -> 2HNO3 + NO


1
Expert's answer
2021-10-07T03:47:57-0400

Molar mass of NO2= 46.0055 g/mol

= 63.5/46.0055= 1.38mol

Mole ratio of NO2: HNO3 = 3:2

= 2×1.38/3 = 0.92mol